Five batteries are connected in parallel to supply a certain load. Each battery consist of 12 cells in series. Each cell has an emf of 2V and the internal resistance of each cell is 0.15Ω. What should be the resistance of the load so that the total power consumed is 384 watts?


1
Expert's answer
2021-11-09T10:44:37-0500

The total power is the total voltage of the battery over the load resistance. The battery voltage:


"V=(12·2)=24\text{ V},"

the total resistance of the battery is


"r=12r/5=0.36\Omega."

The current:


"I=V/(r+R_L),\\
P=VI=V^2/(r+R_L),\\\space\\
R_L=V^2/P-r=1.14\Omega"